OpenStudy (anonymous):

a,b and c are real values satisfying a+2b+3c=195 and a^2+b^2+c^2=2925. The maximum value of a has the form p/q . What is the value of p+q? @LolWolf

OpenStudy (anonymous):

What are you able to use in solving these? I see no elementary way of doing such... maybe I'm missing something entirely, but the only way I can solve this would be using elliptic curves.

OpenStudy (anonymous):

*I can see in solving this

OpenStudy (anonymous):

Just wondering, what are these problems for?

OpenStudy (anonymous):

Hm... Here's where I'm at: \[ 3120-b(b+2)-c(c+3)\equiv0\!\!\mod p\\ \frac{p}{q^2}(p+q)=3120-b(b+2)-c(c+3)\\ \]Where \(p, q\in \mathbb{Z}\) and \(\gcd(p,q)=1 \).

OpenStudy (anonymous):

Sorry to take ur time... Of no use because i'm in class 12 currently.. don't know elliptic curves either...

OpenStudy (anonymous):

It's all right, I need to go, anyways, if you can solve it, awesome; I'd recommend calling the other guys over, though.